Circular No. 8250

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

GRB 031203 P. van Dokkum, Yale University; J. S. Bloom, Harvard-Smithsonian

Center for Astrophysics; and C. Bailyn, B. Cobb, and M. Buxton, Yale University, report further analysis of SMARTS-consortium data on GRB 031203 (cf. GCN 2463) obtained between Dec. 4.125 and 4.258 UT. The summed images reveal no source in I or J (upper magnitude limits I = 23 and J = 21.0) at the Rapid-Eye-Mount (REM) robotic-telescope position from La Silla (cf. GCN 2466). A faint source with J = 20.5 ± 0.5, undetectable in I, is found 2′′.5 from the REM position at α = 8^{h}02^{m}29^{s}.78, δ = −39^{o}51′07′′.1 (equinox 2000.0) ± 0′′.4 (1σ; found from astrometric tie to 2MASS). No other source is found within 4′′.5 of the REM position. The J-band data were obtained starting at Dec. 4.180 (essentially contemporaneous with the REM observations), while the I-band images span the entire three hours of observation. Taking the REM and SMARTS photometry at face value, it is found that the source has the characteristics of a J ‘dropout’, suggesting a redshift of z > 9 (formally z ≥ 11). Models assuming extreme extinction in the host galaxy cannot reproduce the observed spectral en-

ergy distribution (see http://www.astro.yale.edu/dokkum/GRB031203.

html). There are many caveats, and the nature of this source is still highly uncertain. Nevertheless, as this is a potential discovery of the highest redshift source known, the authors encourage the community to obtain additional near-infrared photometry and (particularly) near-infrared spectroscopy of this object.

COMET C/2002 O7 (LINEAR) M. Mattiazzo, Wallaroo, South Australia, reports that a CCD image

taken by himself on Sept. 27.8 UT showed that this comet had apparently disrupted, revealing only a diffuse sunward-pointing ‘antitail’ of debris. Requests from the Central Bureau were sent to several observers at the time, seeking confirmation and additional details. Nothing further was received. But now, G. P. Tozzi (Istituto Nazionale di Astrofisica, Arcetri), along with H. Boehnhardt, O. R. Hainaut, F. Selman, I. Saviane, M. Pizarro, G. Roman, and F. Labrana, report that they searched for the comet on Dec. 3 (when its predicted total magnitude was ∼ 10) with several telescopes at the European Southern Observatory, including the 2.2-m reflector (R filter), the New Technology Telescope (+ JHK filters), and the 3.6-m reflector (+ VRi filters); the comet could not be detected to a limiting magnitude of R ∼ 20.5 within an area 30′ × 30′ centered at its predicted position.

2003 December 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green